OpenStudy (anonymous):

/evaluate cos(alpha+beta) when sin(alpha)= -7/25 and cot(beta)=-8/15; Neither alpha nor beta are in quadrant IV

OpenStudy (anonymous):

 \[\cos \left( \alpha+\beta \right)=\cos \alpha \cos \beta-\]